OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3.

OpenStudy (anonymous):

y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (saifoo.khan):

C part.